find the dimensions of a rectangle whose length is 50 inches more than its width and whose perimeter is 500 inches

Answer by checkley79(3341)

P=2L+2W
L=W+50
500=2(W+50)+2W
500=2W+100+2W
4W=500-100
4W=400
W=400/4
W=100 INCHES FOR THW WIDTH.
L=100+50=150 INCHES FOR THE LENGTH.
PROOF:
500=2*150+2*100
500=300+200
500=500